OpenStudy (anonymous):

given the polynomial f(x), if (x+5) is a factor, what else must be true? please help

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

@robtobey

OpenStudy (phi):

if f(x) has a factor of (x+5) that means f(x)= (x+5) g(x) where g(x) is another polynomial notice if x= -5, f(-5) = (-5+5) * g(-5) = 0*g(-5) = 0 in other words, we know that f(-5)= 0

OpenStudy (anonymous):

thanks man